OpenStudy (anonymous):

Write a rule for the linear function in the table. x:1,2,3,4 y:f(x):-7,-10,-13,-16 A. f(x) = -1/3x – 4 B. f(x) = 3x + 4 C. f(x) = –3x – 4 D. f(x) = x – 7

OpenStudy (hoblos):

-3-4=-7 -6-4=-10 -9-4=-13 -12-4=-16 so f(x) = –3x – 4